Three 2100 Ω resistors, all in parallel, are connected across a 210-V dc source. Calculate (a) the

total current, (b) the current through each resistor, and (c) the total power dissipated in the resistors.



1
Expert's answer
2022-01-10T09:16:19-0500

Explanations & Calculations

  • Just refer to the formula for resistors in parallel.
  • This will lead you to the answer, "\large \frac{2100\Omega}{3}=\small 700\Omega."


  • The same current flows through each resistor as each one is identical.
  • Use Ohm's law to find the current, "\small I=\large\frac{V}{R}=\frac{210\,V}{700\,\Omega}=\small 0.3\,A"


  • The power dissipated (heat) is given by "\small p=I^2R" for a resistor.
  • You can use this formula for this step.
  • Total power dissipation is 3 times that of one.
  • You know the current through a resistor and their resistance, so you can give it a try.
  • Unit should be Joules(J)